OpenStudy (anonymous):

what is the square root of 3+4i? explain please.

OpenStudy (anonymous):

\[ \sqrt{3+4 i}=(a+i b)\\ 3+ 4 i = a^2 - b^2 + 2 a b i a^2 - b^2 =3\\ 2 a b = 4 \] Solve for a and b

OpenStudy (anonymous):

One of the solution is a=2 and b=1

OpenStudy (anonymous):

2 + i is the answer

OpenStudy (anonymous):

yes it has two answers

OpenStudy (anonymous):

Of course -(2 + i) is a solution

OpenStudy (anonymous):

z^1/n has n different solutions .always!

OpenStudy (anonymous):

Unless z =0.

OpenStudy (anonymous):

yeap